# Multi-lingual Functional Evaluation for Large Language Models

### v1

@@ Abstract
Multi-lingual competence in large language models is often evaluated via static data benchmarks such as Belebele, M-MMLU and M-GSM. However, these evaluations often fail to provide an adequate understanding of the practical performance and robustness of models across multi-lingual settings. In response, we create multi-lingual functional benchmarks – Cross-Lingual Grade School Math Symbolic (CL-GSM Symbolic) and Cross-Lingual Instruction-Following Eval (CL-IFEval)– by translating existing functional benchmark templates from English to five additional languages that span the range of resources available for NLP: French, Spanish, Hindi, Arabic and Yoruba. Our results reveal that some static multi-lingual benchmarks capture functional performance much more closely than others (i.e. across models, there is a 24%, 17% and 18% decrease in performance between M-GSM and CL-GSM Symbolic in English, French and Spanish respectively; similarly there’s a 15 - 24% performance drop across languages between Belebele and CL-IFEval, and only a 0.5% to 3% performance drop between M-MMLU and CL-IFEval). Similarly, we find that model robustness across languages varies significantly, with certain languages (eg. Arabic, English) being the most consistently well performing across evaluation iterations.
@@ 1 Introduction
Figure 1: Description of the functional evaluation paradigm. Unlike with static data benchmarks, in the functional evaluation paradigm, model input prompts are not fixed but generated through a fixed template and a set of variables $X$ (modifiable prompt attributes meant to impact model outputs) and a set of distractors $D$ (modifiable prompt attributes meant to be ignored). The ground truth in this setting is generated through a fixed functional transformation $f(X)$. For instance, the prompt "Sally bought 2 red apples and 3 green apples. How much fruit did Sally buy?" is generated from the fixed template "$\{name\}$ bought $\{n_{1}\}$$\{color_{1}\}$ apples and $\{n_{2}\}$$\{color_{2}\}$ apples. How much fruit did $\{name\}$ buy?". This template involves the variables $X=\{n_{1},n_{2}\}$ and the distractors $D=\{name,color_{1},color_{2}\}$. The correct fixed output function in this case is $f(X)=n_{1}+n_{2}$.
@@
Our experiments reveal that there are significant, though inconsistent disparities between multi-lingual model performance on functional vs static data benchmarks – in the vast majority of cases, the models perform much better on the static data benchmarks than the functional evaluations, though these differences are much larger for certain static data benchmarks than others. The performance gap between languages is comparable in both cases but this varies widely across languages and models. Furthermore, functional benchmark results reveal robustness inconsistencies across languages, with models performing more robustly in certain languages over others for specific prompt templates or question types.
Refer to captionRefer to captionRefer to caption
@@ 2 Related Work
Common multi-lingual data benchmarks such as M-MMLU (Hendrycks et al., 2021; Lai et al., 2023a), FLORES (Goyal et al., 2022), BeleBele (Bandarkar et al., 2024), and XLSum (Hasan et al., 2021) possess known limitations. The use of direct translation for many of these benchmarks has been critiqued by some as being devoid of realistic cultural context Romanou et al. (2024); Singh et al. (2024). Furthermore, research reveals that English language benchmark data contamination might distort reported benchmark performance in English or possibly additional languages (as is the case for MMLU Dodge et al. (2021) and GSM Zhang et al. (2024)).
@@
Cross-Lingual IFEval (Ours) 5 541 10 31 - 163 Instruction following
@@
Multilingual MMLU Lai et al. (2023a) 26 13062 N/A N/A MC knowledge recall
@@ 3 Benchmark Datasets and Templates
To make multi-lingual functional benchmarks, we translate 541 English IFEval prompts, each containing at least one verifiable instruction, as well as 5000 question-answer pairs generated from 100 templated symbolic problem variants in the English GSM-Symbolic dataset, into French, Spanish, Yoruba, and Hindi. In line with Bang et al. (2023); Lai et al. (2023b), we refer to languages by their ISO 639-1 language code abbreviations and classify languages by the data ratio of the representation of that language in CommonCrawl (see 3) – these languages were thus selected to cover a spectrum of high-resource and lower-resource contexts. As with prior work Montariol et al. (2022); Lai et al. (2023a); Chen et al. (2023), translations are conducted with Google Translate (Wu et al., 2016). To confirm translation quality, we conduct spot checks (i.e. the manual analysis of 10-20 examples per datasets) with native and proficient language speakers to identify inconsistencies in translation accuracy. We name the resulting functional evaluation datasets Cross-Lingual IFEval (CL-IFEval)and Cross-Lingual GSM Symbolic (CL-GSMSym). All translated prompts can be found in our uploaded supplementary materials. For comparison with a set of multi-lingual static data benchmarks, we use Multilingual MMLU (M-MMLU) Lai et al. (2023a); Hendrycks et al. (2021), Multilingual Grade School Math benchmark (MGSM) Shi et al. (2022), and the Belebele benchmark  Bandarkar et al. (2024). The Cross-Lingual GSM Symbolic and Cross-Lingual IFEval datasets are publicly available.  [^1] https://huggingface.co/datasets/vojewale/Cross-lingualGSMSymbolic [^2] https://huggingface.co/datasets/vojewale/Cross-lingualIFEval
@@ 4 Model Evaluation
For reproducible consistency in reporting, our evaluations spanned open-source large language models, including instruction-tuned and multilingual variants - Aya 23-35B, Aya Expanse-32B, Gemma-2-9B-it, Qwen3-8B, Mistral-7B-Instruct-v0.3, Mixtral-8x7B-Instruct-v0.1. For further comparison on CL-IFEval, we also evaluate on the proprietary models GPT-4o-mini and Claude Sonnet 3.5. Using CL-IFEval, we calculate the prompt-level strict/loose accuracy and the instruction-level strict/loose accuracy as defined in the original IFEval paper Zhou et al. (2023). For CL-GSMSym, we evaluate a random sample of 500 translated question-answer pairs per language using the language-specific 8-shot setting for GSM8K evaluation.
@@ 5 Results
Figure 3: Cross-Lingual IFEval Strict Prompt Accuracy.
Figure 5: Cross-Lingual GSMSym Aya-23-35B Template Comparison Across Languages
@@ 5.1 Average Performance Across Languages & Model Ranking
On the other hand, functional benchmarks in natural language, such as CL-IFEval, show a wide range on even English performance (46% - 81%), with other high and medium resource languages such as French (31%- 58%), Spanish (29% - 54%), and Arabic (20% - 49%) being similarly wide ranging in performance. CL-GSMSym shows a similar pattern, with a wide range of model performance in high and medium resource languages (English: 49% - 87%; French: 39% - 78%; Spanish: 34% - 82%; Arabic: 18% - 79%; Hindi: 15% - 72%). Interestingly, in both cases, performance for low resource languages was more consistently low (e.g. in Yoruba, performance on CL-IFEval ranges from 9% to 16%, and performance on CL-GSMSym ranges from 4% to 14%).
For both static benchmarks, the same models are consistently the most performant - typically Qwen3-8B, followed by either Gemma-2-9b-it or Aya-expanse-32B - with the other models not far behind. For functional benchmarks there is a clear discrepancy between a performant class of models (Qwen3-8B, followed by Aya-expanse-32B, followed by Gemma-2-9b-it) and the rest of the models. Unlike with static data benchmarks, this stark performance discrepancy persists, even for low resource languages. For example, on CL-IFEval, the top three models have an English performance range of 70% to 82%, whereas the bottom three models all have an English performance range consistently around 46%. In Yoruba, this performance class discrepancy is still observed, though the performance gap between models is much more narrow. Comparatively, the top three models on the static data benchmark Belebele perform in English at 90% to 93% and the bottom three models perform at 80% to 86%. Across languages, we can see model rankings change significantly when assessed with static vs functional benchmarking. For example, a typically lower-ranked model like Aya-23-35B, outranks Aya-expanse-32B and Qwen3-8B in Yoruba on the static Belebele benchmark but not in Yoruba on the functional CL-IFEval benchmark. Similarly, the Aya-23-35B model outranks Gemma-2-9b-it generally on the static M-GSM, even though it consistently performs worse than Gemma-2-9b-it on the functional CL-GSMSym benchmark, for the same languages.
@@ 5.2 Language Performance Gap
We define the language performance gap to be the difference of the model’s accuracy on its lowest performant language, and the highest performant language. For high- resourced languages, functional evaluations tend to reveal a much larger performance discrepancy than static data benchmarks (see Figure  2). Notably, Aya-expanse-32B, a Cohere model marketed specifically for its multi-lingual capability, has a 5.89% average error gap between high-resourced languages (en, es, fr) across static data benchmarks, but a 23.47% performance gap between those same languages on the functional benchmark CL-IFEval. Similarly, Qwern3-8b has fairly low language performance gaps across high resourced languages on static data benchmarks like M-GSM (14.4%), M-MMLU (4.78%), and Belebele (3.00%), but a large language performance gap on the functional CL-IFEval benchmark (27.36%). Interestingly, for mathematical reasoning, the opposite can be true – for example, the language performance gap across high resource languages is lower for the functional CL-GSMSym (e.g. for Qwern3-8b, 8.6%) than the static data benchmark M-GSM (eg. for Qwern3-8b, 14.4%).
Similarly, when including low resource language settings, CL-IFEval tends to show more optimistic numbers (a higher average performance, a lower performance gap) than the static data benchmark, Belebele. In Appendix B, Tables 4, 5,6, as well as Figure  2, we provide further details on these results.
@@ 5.3 Instruction and Template-Level Robustness Across Languages
To evaluate instruction-level robustness across languages, we use CL-IFEval to compare model behavior on grouped instruction categories. Figures 6, 7, and 8 show success rates for Aya-23-35B, Gemma-2-9B, and Qwen3-8B respectively. These plots reveal notable variation in cross-lingual generalization, with Yoruba performing the worst across most instruction groups and showing complete failure in the “Start/End” category for all models.
To further probe model robustness under controlled prompt variations, we use a subset of CL-GSMSym consisting of 50 samples generated from a fixed set of 10 mathematical question templates. Figures 9, 10, and 11 display the per-template accuracy of the same three models across languages. Template 3(Appendix H), which involves probabilistic inference, consistently produced the lowest performance across models and languages highlighting a persistent weakness in generalizing probabilistic reasoning across linguistic contexts. Interestingly, performance in certain middle-resourced languages (ar, hi) is just as if not more robust for certain types of problems than high resourced languages (e.g. in CL-GSM Symbolic, see Template 2, 4 and 10 results in Figure  9).
@@ 6 Conclusion
Another limitation is the emphasis in our analysis on open-weight models. While we evaluate gpt-4o-mini and Claude Sonnet 3.5 with CL-IFEval as a commercial baseline, our primary focus remains on open-weight models such as Mixtral-8x7B, Mistral-7B, Gemma2-9B-it, Qwen3-8B and AYA models. This creates a possible inherent bias in the scope of our comparisons, as proprietary models may perform significantly better than their open-weight counter-parts. On the other hand, the consistency and transparency of open-weight models make them the preferable object of study – the incorporation of proprietary models can make results hard to reproduce reliably.
@@ Appendix A Additional Experimental Details
Cross-Lingual IFEval (Ours) 5 541 10 31 - 163 Instruction following
@@
Multilingual MMLU Lai et al. (2023a) 26 13062 N/A N/A MC knowledge recall
@@ A.1 Models
For comparison on CL-IFEval, we also evaluate on the proprietary models GPT-4o-mini and Claude Sonnet 3.5.
@@ A.2 Language Resource Classification
Table 3: Classification of languages based on resource availability in CommonCrawl (CC-MAIN-2025-05).
@@ B.1 High-Resource Language Performance Gaps
Aya-23-35B 15.90 4.4 8.0 4.63 5.00
Mixtral-8x7B-Instruct 16.63 7.0 12.0 3.63 6.89
Mistral-7B-Instruct 12.39 14.0 8.0 7.69 15.88
Aya-expanse-32B 23.47 6.2 9.6 3.84 4.22
Gemma-2-9b-it 21.45 27.4 29.2 6.39 3.66
Qwen3-8b 27.36 8.6 14.4 4.78 3.00
@@ B.2 High to Medium-Resource Language Performance Gaps
Aya-23-35B 15.90 17.6 18.08 22.45
Mixtral-8x7B-Instruct 26.06 33.2 30.65 36.78
Mistral-7B-Instruct 28.48 33.8 27.92 38.44
Aya-expanse-32B 28.84 11.0 15.29 15.45
Gemma-2-9b-it 27.18 27.4 18.20 19.55
Qwen3-8b 34.57 27.4 20.03 23.45

### v2

@@ Abstract
Multi-lingual competence in large language models is often evaluated via static data benchmarks such as Belebele, M-MMLU and M-GSM. However, these evaluations often fail to provide an adequate understanding of the practical performance and robustness of models across multi-lingual settings. In response, we create multi-lingual functional benchmarks – Cross-Lingual Grade School Math Symbolic (CL-GSM Symbolic) and Cross-Lingual Instruction-Following Eval (CL-IFEval)– by translating existing functional benchmark templates from English to five additional languages that span the range of resources available for NLP: French, Spanish, Hindi, Arabic and Yoruba. Our results show that the gap between static and functional evaluations is highly uneven: across models, performance drops from M-GSM to CL-GSM Symbolic by 24%, 17%, and 18% in English, French, and Spanish, while the drop from Belebele to CL-IFEval ranges from 15% to 24% across languages, and the drop from M-MMLU to CL-IFEval is much smaller (0.5% to 3%).Similarly, we find that model robustness across languages varies significantly, with certain languages (eg. Arabic, English) being the most consistently well performing across evaluation iterations.
@@ 1 Introduction
Figure 1: Description of the functional evaluation paradigm. Unlike with static data benchmarks, in the functional evaluation paradigm, model input prompts are not fixed but generated through a fixed template and a set of variables $X$ (modifiable prompt attributes meant to impact model outputs) and a set of distractors $D$ (modifiable prompt attributes meant to be ignored). The ground truth in this setting is generated through a fixed functional transformation $f(X)$. For instance, the prompt "Sally bought 2 red apples and 3 green apples. How much fruit did Sally buy?" is generated from the fixed template $\{name\}$ bought $\{n_{1}\}$ $\{color_{1}\}$ apples and $\{n_{2}\}$ $\{color_{2}\}$ apples. How much fruit did $\{name\}$ buy?". This template involves the variables $X=\{n_{1},n_{2}\}$ and the distractors $D=\{name,color_{1},color_{2}\}$. The correct fixed output function in this case is $f(X)=n_{1}+n_{2}$.
@@
Our experiments show that the relationship between functional and static benchmark performance varies significantly across models, languages, and benchmarks. While models often score higher on static benchmarks than on functional evaluations, the size of this gap varies substantially, with some static benchmarks aligning much more closely with functional performance than others. Language performance gaps also vary significantly across settings as their magnitude and direction shift across specific languages and models. Finally, functional benchmarks surface robustness that varies significantly by language and by prompt structure, revealing that models can be stable for certain languages yet brittle for others on particular instruction types, templates, or question families.
@@ 2 Related Work
Common multi-lingual data benchmarks such as M-MMLU (Hendrycks et al., 2021; Lai et al., 2023b), FLORES (Goyal et al., 2022), BeleBele (Bandarkar et al., 2024), and XLSum (Hasan et al., 2021) possess known limitations. The use of direct translation for many of these benchmarks has been critiqued by some as being devoid of realistic cultural context Romanou et al. (2024); Singh et al. (2024). Furthermore, research reveals that English language benchmark data contamination might distort reported benchmark performance in English or possibly additional languages (as is the case for MMLU Dodge et al. (2021) and GSM Zhang et al. (2024)).
@@
Cross-Lingual IFEval (Ours) 5 371 7 31 - 163 Instruction following
@@
Multilingual MMLU Lai et al. (2023b) 26 13062 N/A N/A MC knowledge recall
@@ 3 Benchmark Datasets and Templates
We construct multi-lingual functional benchmarks by translating (a) the 100 English GSM-Symbolic templates (Mirzadeh et al., 2024) and (b) the 541 English IFEval prompts (Zhou et al., 2023) each containing at least one verifiable instruction into French, Spanish, Yoruba, Hindi, and Arabic. Following Bang et al. (2023); Lai et al. (2023a), we refer to languages by ISO 639-1 codes and select them to span resource levels based on CommonCrawl coverage (Table 3).
Initial translations are produced with Google Translate (Wu et al., 2016). We then conduct native-speaker validation of all translated GSM-Symbolic templates and all translated IFEval prompts before any GSM-Symbolic item instantiation.
Validators and recruitment.
For each language, we recruited at least two native speakers volunteers to validate translations. Validators were recruited primarily from students studying at a US-based university. For Yoruba, we additionally recruited a native speaker based in Nigeria to ensure dialectal and cultural appropriateness in a low-resource setting.
Quality control, rubric, and prompt filtering.
Validation followed a structured rubric (Appendix G). Validators compared each translated prompt or template against the English source to ensure fidelity to meaning, intent, and constraint strength, and to verify that instruction realizations remained valid in the target language. Validators discussed and, when possible, corrected problematic translations, and also checked to detect misspellings, duplicated sentences, and template–instance alignment issues. Validators noted that occasional literal word choices remained but did not alter functional constraints or ground-truth mappings.
As part of this validation, they also flagged a subset of original IFEval prompts whose instructions are not meaningfully evaluable in scripts without case distinction (e.g., uppercase/lowercase constraints in Arabic and Hindi). We removed these prompts to preserve cross-lingual comparability and evaluation validity, yielding a 371-prompt CL-IFEval set.
After validation, each GSM-Symbolic template was instantiated into 50 variants per language, yielding 5,000 QA pairs per language, while the validated IFEval prompts constitute CL-IFEval.
For comparison with multi-lingual static benchmarks, we evaluate on M-MMLU (Lai et al., 2023b; Hendrycks et al., 2021), M-GSM (Shi et al., 2022), and Belebele (Bandarkar et al., 2024). The Cross-Lingual GSM Symbolic and Cross-Lingual IFEval datasets are publicly available.  [^1] https://huggingface.co/datasets/vojewale/Cross-lingualGSMSymbolic [^2] https://huggingface.co/datasets/vojewale/Cross-lingualIFEval
@@ 4 Model Evaluation
For tasks that benefit from step-by-step reasoning, we use Chain-of-Thought (CoT) prompting in the same language as the query where applicable. Unless otherwise stated, CL-GSMSym follows the native 8-shot GSM8K setting, and MGSM uses the native_cot configuration. All evaluations are orchestrated with the EleutherAI LM Evaluation Harness (lm-eval), which standardizes prompts, decoding, and metrics and is widely used for open-weight models (Gao et al., 2024).
For reproducible consistency in reporting, our evaluations spanned open-source large language models, including instruction-tuned and multilingual variants—Aya 23-35B, Aya Expanse-32B, Gemma-2-9B-it, Qwen3-8B, Mistral-7B-Instruct-v0.3, and Mixtral-8x7B-Instruct-v0.1. Using CL-IFEval, we compute prompt-level strict/loose and instruction-level strict/loose accuracy as in Zhou et al. (2023). For CL-GSMSym, after template translation, we instantiate 50 variants per template per language and evaluate a random sample of 500 items per language under the native 8-shot GSM8K setting.
Refer to captionRefer to captionRefer to caption
@@ 5.1 Average Performance Across Languages & Model Ranking
Functional benchmarks expose larger within-language variability.
On the other hand, functional benchmarks in natural language, such as CL-IFEval, show a wide range on even English performance (56% - 88%), with other high and medium resource languages such as French (36%- 62%), Spanish (33% - 57%), and Arabic (20% - 50%) being similarly wide ranging in performance. CL-GSMSym shows a similar pattern, with a wide range of model performance in high and medium resource languages (English: 49% - 87%; French: 39% - 78%; Spanish: 34% - 82%; Arabic: 18% - 79%; Hindi: 15% - 72%). Interestingly, in both cases, performance for low resource languages was more consistently low (e.g. in Yoruba, performance on CL-IFEval ranges from 12% to 23%, and performance on CL-GSMSym ranges from 4% to 14%).
Model tiering and ranking shifts under functional evaluation.
For both static benchmarks, the same models are consistently the most performant - typically Qwen3-8B, followed by either Gemma-2-9b-it or Aya-expanse-32B - with the other models not far behind. For functional benchmarks there is a clear discrepancy between a performant class of models (Qwen3-8B, followed by Aya-expanse-32B, followed by Gemma-2-9b-it) and the rest of the models. Unlike with static data benchmarks, this stark performance discrepancy persists, even for low resource languages. For example, on CL-IFEval, the top three models have an English performance range of 70% to 88%, whereas the bottom three models all have an English performance range consistently around 60%. In Yoruba, this performance class discrepancy is still observed, though the performance gap between models is much more narrow. Comparatively, the top three models on the static data benchmark Belebele perform in English at 90% to 93% and the bottom three models perform at 80% to 86%. Across languages, we can see model rankings change significantly when assessed with static vs functional benchmarking. For example, a typically lower-ranked model like Aya-23-35B, outranks Aya-expanse-32B and Qwen3-8B in Yoruba on the static Belebele benchmark but not in Yoruba on the functional CL-IFEval benchmark. Similarly, the Aya-23-35B model outranks Gemma-2-9b-it generally on the static M-GSM, even though it consistently performs worse than Gemma-2-9b-it on the functional CL-GSMSym benchmark, for the same languages.
@@ 5.2 Language Performance Gap
We define the language performance gap to be the difference of the model’s accuracy on its lowest performant language, and the highest performant language. For high- resourced languages, functional evaluations tend to reveal a much larger performance discrepancy than static data benchmarks (see Figure  2). Notably, Aya-expanse-32B, a Cohere model marketed specifically for its multi-lingual capability, has a 5.89% average error gap between high-resourced languages (en, es, fr) across static data benchmarks, but a 25.88% performance gap between those same languages on the functional benchmark CL-IFEval. Similarly, Qwen3-8b has fairly low language performance gaps across high resourced languages on static data benchmarks like M-GSM (14.4%), M-MMLU (4.78%), and Belebele (3.00%), but a large language performance gap on the functional CL-IFEval benchmark (31%). Interestingly, for mathematical reasoning, the opposite can be true – for example, the language performance gap across high resource languages is lower for the functional CL-GSMSym (e.g. for Qwern3-8b, 8.6%) than the static data benchmark M-GSM (eg. for Qwern3-8b, 14.4%). Similarly, when including low resource language settings, CL-IFEval tends to show more optimistic numbers (a higher average performance, a lower performance gap) than the static data benchmark, Belebele. In Appendix B, Tables 4, 5,6, as well as Figure  2, we provide further details on these results.
@@ 5.3 Instruction and Template-Level Robustness Across Languages
Instruction-group breakdowns reveal where cross-lingual generalization fails.
To evaluate instruction-level robustness across languages, we use CL-IFEval to compare model behavior on the IFEvalZhou et al. (2023) default grouped instruction categories. Figures 3, 4, and 5 show success rates for Aya-23-35B, Gemma-2-9B, and Qwen3-8B respectively. These plots reveal notable variation in cross-lingual generalization, with Yoruba performing the worst across most instruction groups and showing complete failure in the “Start/End” category for all models.
Template-controlled math variants isolate which problem families drive errors.
To further probe model robustness under controlled prompt variations, we use a subset of CL-GSMSym consisting of 50 samples generated from a fixed set of 10 mathematical question templates from the GSM Symbolic datasetMirzadeh et al. (2024). Figures 6, 7, and 8 display the per-template accuracy of the same three models across languages. Template 3(Appendix I), which involves probabilistic inference, consistently produced the lowest performance across models and languages highlighting a persistent weakness in generalizing probabilistic reasoning across linguistic contexts. Interestingly, performance in certain middle-resourced languages (ar, hi) is just as if not more robust for certain types of problems than high resourced languages (e.g. in CL-GSM Symbolic, see Template 2, 4 and 10 results in Figure  6).
@@ 6 Conclusion
Another limitation is the emphasis in our analysis on open-weight models. Our primary focus remains on open-weight models such as Mixtral-8x7B, Mistral-7B, Gemma2-9B-it, Qwen3-8B and AYA models. This creates a possible inherent bias in the scope of our comparisons, as proprietary models may perform significantly better than their open-weight counter-parts. On the other hand, the consistency and transparency of open-weight models make them the preferable object of study – the incorporation of proprietary models can make results hard to reproduce reliably.
@@ Appendix A Additional Experimental Details
Cross-Lingual IFEval (Ours) 5 371 7 31 - 163 Instruction following
@@
Multilingual MMLU Lai et al. (2023b) 26 13062 N/A N/A MC knowledge recall
@@ A.2 Language Resource Classification
Table 3: Classification of languages based on resource availability in CommonCrawl ($C$-MAIN-2025-05).
@@ B.1 High-Resource Language Performance Gaps
Aya-23-35B 26.68 4.4 8.0 4.63 5.00
Mixtral-8x7B-Instruct 22.36 7.0 12.0 3.63 6.89
Mistral-7B-Instruct 18.60 14.0 8.0 7.69 15.88
Aya-expanse-32B 25.88 6.2 9.6 3.84 4.22
Gemma-2-9b-it 24.80 27.4 29.2 6.39 3.66
Qwen3-8b 31.00 8.6 14.4 4.78 3.00
@@ B.2 High to Medium-Resource Language Performance Gaps
Aya-23-35B 26.68 17.6 18.08 22.45
Mixtral-8x7B-Instruct 31.41 33.2 30.65 36.78
Mistral-7B-Instruct 37.74 33.8 27.92 38.44
Aya-expanse-32B 28.03 11.0 15.29 15.45
Gemma-2-9b-it 28.04 27.4 18.20 19.55
Qwen3-8b 37.73 27.4 20.03 23.45